OpenStudy (anonymous):

HELP! find the exact value of the real number y in the equation Y=cos-1(√2/2) and explain how you got this answer please?

OpenStudy (jdoe0001):

are you doing the 30-30-60 rule? or doing trig?

OpenStudy (anonymous):

Will you show me both ways? There wasn't a specific way we had to solve it, we just had to solve it.

OpenStudy (jdoe0001):

yes, but before you have to "solve" it, you also need to have it "covered"

OpenStudy (jdoe0001):

so, thus I dunno what section you're doing, for I'd assume your textbook has such

OpenStudy (anonymous):

How do I cover it? I have the unit circle and the text book.

OpenStudy (jdoe0001):

.. "cover" as in, you already had a class on it

OpenStudy (jdoe0001):

I mean, what if the teacher decides to ask you how many planets in andromeda or something, I mean, if it isn't covered... there's no point in testing on it

OpenStudy (anonymous):

Oh she has covered it. I just don't understand trig.

jimthompson5910 (jim_thompson5910):

look at the unit circle and look for points where the x coordinate is sqrt(2)/2 you'll notice that the angle at this point is theta = 45 degrees or pi/4 radians so depending on which mode you're in, you'll get Y=cos-1(√2/2) Y=45 degrees ... if you want an answer in degrees or y = pi/4 radians ... if you want an answer in radians

OpenStudy (anonymous):

Oh that is actually pretty simple. Thank you :)

jimthompson5910 (jim_thompson5910):

yw